deathsindragon:

a hat salesman earns a commission of 30% of all hat sales made. yesterday he sold 3 hats for $90 each and 4 hats for $50 each. how much did he earn in commissions yesterday? A. $60 B. $141 C. $81 D. $42

XioGonz:

Your answer is B and I'll explain why. The percentage of commission that was earned by the hat salesman is = 30 percent. The amount of money for which 3 of the hats were sold is= 90 * 3 So the outcome = 270 dollar And the amount of money for which 4 hats were sold = 50 * 4 And that outcome = 200 dollars Then the total amount of money for which 7 hats were sold = 200 + 270 That'll be = 470 dollars The amount of commission that was earned by the hat salesman = 470 * (30/100) So your answer is = 141 dollars The amount of commission earned by the hat salesman was $141.

XioGonz:

I get best answer?

deathsindragon:

yes u do xio, thx

XioGonz:

Thank youu